Please help me with this question: Name the x-intercepts and the vertex of the graph of y=(x+3)(x-1). Then sketch the graph of the function.
Answer by stanbon(57384)

To find the x-intercepts, let y=0 and solve for "x", as follows:
0=(x+3)(x-1)
Since the product is zero one of the factors is zero.
If x+3=0 then x=-3
If x-1=0 then x=1
x-intercepts are (-3,0) and (1,0)
VERTEX:
Rewrite the equation as follows:
y=x^2+2x-3
Complete the square, as follows:
y+3=x^2+2x
y+3+1=x^2+2x+1
y+4=(x+1)^2
Vertex is at x=-1, y=-4
Cheers,
Stan H.